OpenStudy (anonymous):

Solve this equation over the interval [0,2π) tan^2 x -3=0

OpenStudy (tkhunny):

Can you solve this for z? z^2 - 3 = 0

OpenStudy (anonymous):

Z^2-3=0 +3=+3 z^2=3

OpenStudy (tkhunny):

Why did you stop? Keep going. Make sure you find BOTH results. When you are done, do EXACTLY the same thing with your original equation.